# Hallucination Detection in Foundation Models for Decision-Making: A Flexible Definition and Review of the State of the Art

### v1

@@ 4.1.1 Hidden States
4.1.2 Attention Weights
  1. 1.
    OPERA, proposed by Huang et al. (2024a), is a hallucination detection method for LVLMs that makes use of the model’s internal attention weights. When visualizing the attention matrix, the authors find that there exist peculiar column patterns that align with the beginning of a hallucinated phrase. These aggregation patterns usually occur on a non-substantial token like a period or quotation mark, but are deemed to have a large impact on the prediction of future tokens. As such, this finding led Huang et al. (2024a) to modify the beam search algorithm Freitag and Al-Onaizan (2017) by applying a penalty term to beams wherever an aggregation pattern is detected, and roll back the search to before the pattern arises. Their method is shown to reduce hallucinations, and even eliminate possible repetitions in generations.
4.1.3 Honesty Alignment
  1. 1.
    For example, Lin et al. (2022a) collect a calibration dataset of questions and answers from GPT-3 under multiple types of tasks (e.g., add/subtract and multiply/divide), and record how often each task is incorrectly answered. They aim to fine-tune the LLM to also output its certainty that the prediction is correct. Consequently, Lin et al. (2022a) fine-tune the model with data pairs of a question and the empirical accuracy on the task that the question originates from in the calibration dataset, such that the model is expected to similarly output a probability of accuracy at test-time. The authors show that the proposed verbalized probability in deployment does correlate with actual accuracy on the tasks.
  2. 2.
    Yang et al. (2023a) take the method one step further by also training the model to refuse to answer questions with high uncertainty.
4.2 Grey-box Methods
4.2.1 Concept Probabilities
  1. 1.
    Empirically, Varshney et al. (2023) show that there is a negative correlation between hallucination rate and token probability (i.e., as a token’s probability decreases within a sentence, the tendency to hallucinate increases). Thus, the authors rely on token probabilities to estimate uncertainty of concepts within a generated claim, and they check for correctness by cross-referencing a knowledge-base. Whenever a concept is found to be conflicting with a fact through verification questions, their method attempts to mitigate the error by prompting the LLM to replace the incorrect claim with the evidence. Although effective in the QA setting, Varshney et al. (2023) concede that, in the event token probabilities are not available, some form of heuristic must be used to detect hallucination candidates.
  2. 2.
  3. 3.
    SayCanPay, proposed by Hazra et al. (2024), builds off of the SayCan framework Ichter et al. (2023) to improve the expected payoff of following a plan specified by a language model. Within our hallucination definition, this goal translates to increasing the desirability of generations by improving the likelihood of the model achieving higher rewards. The authors propose three different strategies for planning: Say, SayCan, and SayCanPay. Say methods greedily choose next actions based only on token probabilities. SayCan approaches also take the success rate of the chosen action into consideration. Finally, SayCanPay additionally estimates the expected payoff from following the plan with some heuristic. Hazra et al. (2024) learn this Pay model with regression on an expert trajectory dataset. Combining all three models together minimizes the likelihood that a generated plan contains conflicting infeasible action calls, while maximizing the efficiency of the task completion.
4.2.2 Conformal Prediction
Another range of works estimate the uncertainty of a model output with conformal prediction so as to provide statistical guarantees on the likelihood of predictions being correct Shafer and Vovk (2008).
  1. 1.
    Quach et al. (2024) propose conformal language modeling to build a set of possible candidate responses to a test prompt, while calibrating algorithm parameters on a held-out dataset of independent prompts and their corresponding admission functions, which check whether a model output meets the criteria of an input prompt. In their algorithm, the authors calibrate thresholds for three separate scoring functions that test for generation quality, similarity with other responses, and model confidence using “Learn then Test” Angelopoulos et al. (2022). At inference time, given scoring functions, calibrated thresholds, and an input prompt, the method samples outputs from the model and adds them to a prediction set if they meet quality and similarity thresholds, until the whole set is guaranteed to meet the user-defined confidence parameter.
  2. 2.
@@ 4.2.2 Conformal Prediction
  3. 3.
    While the previous hallucination mitigation works presented using conformal prediction are solely applied to QA settings, Ren et al. (2023) are the first to apply conformal prediction of foundation models to robotic tasks. The authors are motivated by a desire for language-conditioned robots to understand when they are uncertain about the next action to take, such that they can ask for help in those cases (while minimizing frequency of clarifications). Because LLM generations with different length sequences inherently produce different complete sentence probabilities, the authors propose framing the control task as a multiple-choice problem, like Kumar et al. (2023). Their approach, KnowNo, prompts an LLM to generate a possible set of next actions to take in multiple choice form. They first collect and hand-label a calibration dataset of pairs of held-out instructions and the probability of the model choosing the best action to take next. At test-time, the model outputs a set of actions to take and KnowNo eliminates actions with token probabilities less than a calibrated certainty from a user. If there are still multiple actions left in the prediction set after eliminating uncertain actions, the model queries a human for help choosing the next action. Ren et al. (2023) show that KnowNo deviates from the user-defined error rate least often compared to methods that do not use conformal prediction and has the highest success-to-clarification ratio. However, several assumptions had to be made to produce the demonstrated results, including having access to next token probabilities, having resources to collect a large calibration dataset, presuming people will faithfully provide help when asked for it, and using ground-truth vision to fully ground the environmental objects with the text input to the model.
  4. 4.
    Liang et al. (2024) extend the KnowNo methodology by incorporating an introspective planning step using a previously constructed knowledge-base of experiences, which tends to (1) enhance quality of generated plans, and (2) improve interpretability of decisions. Specifically, introspective planning first constructs a knowledge-base containing training pairs of tasks, observations, and valid plans, which the LLM is prompted to generate explanations behind why they are reasonable. Each experience is stored with a key as an embedding of the original instruction. During inference, given a new test instruction, their method queries the database to find the key with the closest embedding to that of the new instruction. This previous experience and reasoning is fed into the model to generate a set of candidate plans to follow. Finally, the remainder of the algorithm follows the same process as KnowNo to calibrate and narrow down the prediction set to fall within a desired error rate.
  5. 5.
    Wang et al. (2024) aim to provide additional guarantees on completing the task provided within the natural language instruction. To do so, the authors propose a novel task specification, LTL-NL, which combines linear-temporal-logic (LTL) descriptions with natural language from a user instruction, which the authors claim is easier to define than classical LTL specifications. Given this specification, a symbolic task planner chooses a sub-task to complete next and an LLM generates plans for each sub-task, respectively. Like Ren et al. (2023) and Liang et al. (2024)Wang et al. (2024) apply conformal prediction to minimize the number possible actions to take next within some desired error rate. However, rather than directly asking a user for assistance when there is high uncertainty in the next action to take (or when there are environmental constraints), their method, HERACLEs, samples a new sub-task to complete from the task planner. If on the other hand, the task planner is unable to provide a new sub-task, HERACLEs requests help from the user. With experimentation, the authors find that their method achieves higher task completion rate on missions requiring more sub-tasks, outperforming baseline planners that do not utilize LTL specifications.
4.3 Black-box Methods
4.3.1 Analyzing Samples from Model
  1. 1.
    For example, SelfCheckGPT, proposed by Manakul et al. (2023), samples multiple responses (each of which may contain many sentences) from an LLM to a single query, and measures the consistency among the varied responses through a study with five different approaches. SelfCheckGPT with BERTScore Zhang et al. (2020), for example, computes a similarity score between two sentences from different sampled outputs. Intuitively, a hallucination is detected when the similarity score for a sentence is low across all other samples. Other consistency-checking methods the authors consider include using an automatic multiple-choice QA system conditioned on the responses, relying on a proxy-LLM which has access to token probabilities, training an external classifier to predict contradictions (coined SelfCheck-NLI), and directly prompting the LLM to evaluate whether any sentence can be supported by another sampled response. As expected, Manakul et al. (2023) find that sampling more responses from the model leads to better estimation of the validity of a claim, but is slower to compute.
  2. 2.
@@ 4.3.1 Analyzing Samples from Model
    A concurrent work from Elaraby et al. (2023) rather computes the entailment among responses at the sentence-level. Consistency metrics check whether responses contradict one another while entailment metrics identify if the responses imply one another. The nuanced difference between SelfCheck-NLI and their method, HaloCheck, is that Elaraby et al. (2023) uses the SummaC Laban et al. (2022) entailment estimation method, placing equal weightage among all sentences, in all responses, to compute a more balanced prediction score. HaloCheck is also shown to be more efficient at predicting scores. Other commonly used metrics within the language community for similarity estimation are discussed in Section 5.1.1.
  3. 3.
    Rather than analyzing the responses of a single language model, Du et al. (2023) take an ensemble approach. Specifically, they propose pitting multiple instances of an LLM into a debate of the correct answer to a question. In practice, several agents are given the same question and predict a response. Over multiple iterations, all the responses of other agents are concatenated, and fed in as additional context to each model, and a new response is sampled. The authors show that a combination of their iterative ensemble approach with chain-of-thought reasoning mitigates individual hallucinations (as agents tend to converge on a single consensus) and increases QA accuracy. While only tested on offline datasets, the approach could be utilized within a simulation framework, like the one presented by Park et al. (2023), where, for example, multiple language agents may debate about plans to make.
  4. 4.
    CLARA is a framework engineered by Park et al. (2024) that predicts when an instruction provided to a robotic system controlled by an LLM may be ambiguous or infeasible. Intuitively, if a language model is uncertain about an instruction, it might output diverse (or conflicting) actions to other instructions that hold similar information. Thus, CLARA samples several sets of concepts from the original prompt, randomly orders them to assemble multiple inputs to the model, and passes them as input to the language model under test. The method computes the average similarity of pairs of output actions in an embedding space, over all outputs. Next, to check for infeasibility of the original goal, the foundation model is provided the possible action space of the robot, environmental observation, and goal, and is prompted to output whether the desired task is practical. In the event the model is uncertain from the multi-prompting step, but the goal is feasible, CLARA asks for clarification from the user with reasoning for why it is uncertain. While the method achieves a reasonable success rate on robotic pick-and-place tasks with real user instructions (where other discussed methods are primarily evaluated in QA settings), there are still failure cases where the model hallucinates during uncertainty reasoning and feasibility prediction.
  5. 5.
    Another set of works explicitly identify contradictions among responses, instead of estimating similarity. Naturally, detecting a self-contradiction is guaranteed to reveal an invalid claim. Mündler et al. (2024) pose that removing detected conflicting information will increase the validity of a generated response. As such, the authors suggest a solution that finds important concepts within a response to be evaluated, prompts the generation model to generate more information about each of the concepts, and uses a separate analyzer language model to evaluate the consistency of pairs of sentences on the same concept. Any sentences that are found to be conflicting are revised by the analyzer model, before being output to the user.
  6. 6.
    Rather than relying on another language model to analyze the correctness of predictions, Dhuliawala et al. (2023) utilize chain-of-thought reasoning to prompt an LLM to generate possible verification questions about its original responses. If there are conflicts in answers to the verification questions, the LLM is prompted to regenerate its output with updated context and conflicting reasoning.
  7. 7.
@@
    In the problem space of object detection, Li et al. (2023c) estimate the accuracy of a (possibly hallucinating) LVLM when describing an image with a text caption. Early metrics, like CHAIR Rohrbach et al. (2018), fail to provide stable estimates of accuracy when different captions with similar semantic grounding lead to varying scores. To tackle this stability problem, the authors propose POPE, which curates binary questions about whether an object exists within an image scene. Questions to which the foundation model provides conflicting responses describe objects that the model may be hallucinating. These metrics are further described in Section 5.1.2.
  8. 8.
@@
    Yet another recent work by Xiong et al. (2024) asks LLMs in a zero-shot manner to verbally include their uncertainty in the generated output. This desired behavior is elicited through additional prompt engineering (i.e., add a phrase to the prompt like “Please provide your confidence level as a percentage”). Unlike Lin et al. (2022a), the authors do not further fine-tune the model to a calibration dataset of uncertainties. To combat over-confidence in output scores, Xiong et al. (2024) utilize chain-of-thought reasoning, predict the confidence score of each sub-claim, and combine them over the whole response to compute the final belief. The authors find that a hybrid approach, merging verbalized uncertainty with self-contradiction detection, performs better than the individual components alone on expected calibration error, comparing predicted confidence and actual model accuracy.
4.3.2 Adversarial Prompting
  1. 1.
    For example, Mehrabi et al. (2023) apply adversarial prompting to text-to-image foundation models, like Stable Diffusion Schramowski et al. (2023), to generate offensive images. With respect to Definition 3.1, their framework, FLIRT, is essentially testing the tendency of foundation models to hallucinate undesired generations in deployment. FLIRT uses an adversarial language model to predict a prompt to input to the image generator, scores the generated image for the presence of undesirable traits using an external classifier, re-prompts the adversary to produce a new instruction conditioned on the findings of the classifier, and repeatedly generates images until the adversary successfully prompts the test model to output an undesirable result. Mehrabi et al. (2023) define objective functions conditioned on the score output by external classifiers to maximize diversity of adversarial prompts and minimize toxicity so as to pass text filters that detect malicious inputs, while improving attack effectiveness.
  2. 2.
    Another work from Yu et al. (2023) presents the AutoDebug framework for automatically sampling and updating several prompts for use in adversarial testing of the language model. The authors specifically explore adversarial testing under the case that the model predicts a correct response when provided relevant context, but generates an incorrect prediction when the evidence is modified. They apply two different modification approaches: replacing tokens within the context to provide incorrect facts, and adding additional relevant facts to the prompt that may make it difficult to pick out the most important details.
4.3.3 Proxy Model
  1. 1.
    One such method is used as a baseline within the SelfCheckGPT article Manakul et al. (2023). As many language foundation models do not provide access to token probabilities, the authors use an open-source proxy LLM that does provide token probabilities as an estimate of the original output’s probability. They find that using proxy LLMs for probability estimation and hallucination detection successfully is highly variable. The accuracy of detection is dependent on the complexity of the LLM itself, as well as the training data of the proxy LLM (i.e., models trained on independent datasets from the original LLM will have different generation patterns).
  2. 2.
    Within this section, we also include works that use an external trained classifier to detect hallucinations. For example, Chen et al. (2023d) curate a dataset of QA dialogue from LLM generated responses. They apply a composition of metrics to assess quality of responses, including a self-assessment from the LLM comparing the ground-truth and predicted text, human-labeled, and machine metrics (e.g., BERT score, F1 score, BLEU, etc.). Their hallucination discriminator, RelD, is trained on the dataset in multiple separate phases, each using a different objective: regression, multi-class classification, and finally binary classification. Through experiments, they find that RelD closely aligns with human evaluators’ original predictions.
  3. 3.
@@ 4.3.3 Proxy Model
    Similarly, Pacchiardi et al. (2024) develop a black-box lie detector for LLMs. In their case, the authors hypothesize that models that output a lie will produce different behaviors in future responses, like Azaria and Mitchell (2023). As such, at inference time, Pacchiardi et al. (2024) prompt the LLM with several binary questions (that may be completely unrelated to the original response) and collect yes/no answers. All the responses are concatenated into a single embedding that is input to the logistic regression model to predict the likelihood that the response was untruthful. The authors find that the simple detector is mostly task- and model-agnostic once trained on a single dataset.
4.3.4 Grounding Knowledge
  1. 1.
    PURR, proposed by Chen et al. (2023a), is a denoising agent, like LURE (discussed in Section 4.2.1), that is trained in an unsupervised fashion given evidence from online sources, a clean (correct) summary, and a noisy (hallucinated) summary. The model learns to denoise the incorrect summary to the clean statement. During deployment, given a possibly hallucinated claim, a question generation model queries online sources for evidence about the claim, and PURR generates a cleaned version of the original summary with said evidence.
  2. 2.
@@ 4.3.4 Grounding Knowledge
  3. 3.
    Another source of potential conflict that leads to hallucinations, is misalignment between a model’s capabilities and the user’s beliefs about what it can do. Zhang et al. (2023c) tackle this knowledge alignment problem and categorize alignment failures into four types:
@@
  4. 4.
    Peng et al. (2023) aim to add plug-and-play modules to an LLM to make its outputs more accurate, since these large foundation models cannot feasibly be fine-tuned whenever there is new information. Their work formulates the user conversation system as a Markov decision process (MDP) whose state space is an infinite set of dialogue states which encode the information stored in a memory bank, and whose discrete action space includes actions to call a knowledge consolidator to summarize evidence, to call an LLM prompt engine to generate responses, and to send its response to the user if it passes verification with a utility module. The proposed LLM-Augmenter has a memory storing dialogue history, evidence from the consolidator, set of output responses from an LLM, and utility module results. Its policy is trained in multiple phases with REINFORCE Williams (1992) starting with bootstrapping from a rule-based policy designed from domain experts, then learning from simulators, and finally, from real users. The authors find that access to ground-truth knowledge drastically improves QA results, and feedback from the utility module and knowledge consolidator help to provide more accurate answers to users.
  5. 5.
    Evaluated in actual decision-making settings, Introspective Tips Chen et al. (2023b) provide concise, relevant information to a language planner to learn to solve more efficiently. Intuitively, summaries that collect information over all past experiences may be long and contain unnecessary information. In contrast, tips are compact information with high-level guidance that can be learned from ones own experiences, from other demonstrations, and from other tasks in a similar setting. Chen et al. (2023b) show that providing low-level trajectories is less effective than tips on simulated planning tasks. Additionally, with expert demonstrations, the LLM learns faster with fewer number of failed trials than with just past experience alone. However, one limitation identified in the study is that the LLM underperforms in unseen, low-difficulty missions where it has issues generating general tips for zero-shot testing.
4.3.5 Constraint Satisfaction
There is also additional work in creating black-box algorithms for ensuring decision plans generated by foundation models meet user-defined goal specifications and system constraints, like their grey-box counterpart developed by Wang et al. (2024).
  1. 1.
    Because these models under test provide their results in text form, it is natural to apply formal method approaches (e.g., satisfiability modulo theory, SMT, solvers) to verify the satisfaction of generated plans. For example, Jha et al. (2023) prompt an LLM planner with a problem formulated with first order constraints to predict a set of actions to complete the task. The output plan is input to an SMT solver to check for any infeasibilities in the program, and any counterexamples found are used to iteratively update the prompt and generate new plans. This counterexample approach is much faster than relying on combinatorial search methods that find a plan from scratch. However, the quality of generated plans and the number of iterations before a successful plan is generated are heavily dependent on the LLM generator itself, with similar reasons to the proxy-model used by Manakul et al. (2023).
  2. 2.
    Another work from Hu et al. (2024) develops a RoboEval benchmark to test generated plans on real robots, in a black-box manner. Like Wang et al. (2024), the authors introduce their own extension of LTL formulations, known as RTL, which specifies temporal logic at a higher, scenario-specific, level, while abstracting away constraints that are not dependent on available robot skills. RTL and LTL-NL are easier to read and define than classic LTL methods. RoboEval utilizes the provided RTL formulation of a problem, a simulator, and evaluator to systematically check whether the output meets requested goals. Furthermore, to check for robustness of the model to varied instructions, Hu et al. (2024) hand-engineer paraphrased sentences within an offline dataset that should ideally result in the same task completion. Primary causes of failures were found to be a result of generated code syntax/runtime errors, attempting to execute infeasible actions on the robot, and failing RTL checks.
5 Metrics and Evaluation Platforms

### v2

@@ C.1.1. Hidden States
C.1.3. Honesty Alignment
  1. (2)
    Yang et al. (2024b) take the method one step further by also training the model to refuse to answer questions with high uncertainty.
C.2. Grey-box Methods
C.2.2. Conformal Prediction
  1. (2)
@@ C.2.2. Conformal Prediction
C.3. Black-box Methods
C.3.1. Analyzing Samples from Model
  1. (2)
@@ C.3.1. Analyzing Samples from Model
    A concurrent work from Elaraby et al. (2023) rather computes the entailment among responses at the sentence-level. Consistency metrics check whether responses contradict one another while entailment metrics identify if the responses imply one another. The nuanced difference between SelfCheck-NLI and their method, HaloCheck, is that Elaraby et al. use the SummaC (Laban et al., 2022) entailment estimation method, placing equal weightage among all sentences, in all responses, to compute a more balanced prediction score. The authors evaluate HaloCheck and hallucination mitigation techniques on a domain-specific, custom-curated dataset, with facts about the US National Basketball Association (NBA). Specifically, Elaraby et al. first prompt GPT-4 for questions on topics within the NBA domain and manually filter out low-quality generations, resulting in $151$ questions. An LLM is then queried for $5$ responses to each question, and each response is manually annotated for consistency and accuracy of generations. HaloCheck is also shown to be more efficient at predicting scores. Other commonly used metrics within the language community for similarity estimation are discussed in Appendix D.1.1.
  2. (7)
@@
    In the problem space of image captioning, Li et al. (2023b) estimate the accuracy of a (possibly hallucinating) LVLM when describing an image with a text caption. Early metrics, like CHAIR (Rohrbach et al., 2018), fail to provide stable estimates of accuracy when different captions with similar semantic grounding lead to varying scores. To tackle this stability problem, the authors propose POPE, which curates binary questions about whether an object exists within an image scene. Questions to which the foundation model provides conflicting responses describe objects that the model may be hallucinating. These metrics are further described in Appendix D.1.2.
  3. (8)
@@
    Yet another recent work by Xiong et al. (2024) asks LLMs in a zero-shot manner to verbally include their uncertainty in the generated output. This desired behavior is elicited through additional prompt engineering (i.e., add a phrase to the prompt like “Please provide your confidence level as a percentage”). Unlike Lin et al. (2022a), the authors do not further fine-tune the model to a calibration dataset of uncertainties. To combat over-confidence in output scores, Xiong et al. utilize chain-of-thought reasoning, predict the confidence score of each sub-claim, and combine them over the whole response to compute the final belief. The authors find that a hybrid approach, merging verbalized uncertainty with self-contradiction detection, outperforms the individual components alone on expected calibration error, comparing predicted confidence and actual model accuracy.
  4. (9)
    Yehuda et al. (2024) present InterrogateLLM, a sampling-based approach that attempts to reconstruct an original query given a possibly hallucinated LLM response, and measures the similarity of the two queries. Inspired by human studies (Brewer et al., 1999), the authors specifically hypothesize that responses that generate queries that differ greatly from the original prompt point to possible hallucinations. However, they also confess that, like human studies, there is the possibility of false positive hallucination detections due to the stochastic nature of language models. InterrogateLLM follows a simple procedure: form a prompt with few-shot examples containing queries and answers followed lastly by the actual query, generate a (hallucinated) response from a language model under test, reverse the original prompt examples to provide answers and their queries with the generated response appended last, sample generated queries from any language model (not necessarily the one under test), and measure the cosine similarity of vector embeddings between the original query and generated queries. Using public datasets covering a range of trivia knowledge (Banik, 2017; Ziegler, 2020; Elgiriyewithana, 2023), the authors curate a dataset of QA pairs. Most importantly, the authors find that their backward few-shot method is critical to decreasing the false-negative detection rate of hallucinations seen in SelfCheckGPT (Manakul et al., 2023), which only compares generated responses from the original forward pass. Furthermore, using an ensemble of models for the backward generation and more iterations of query sample generation lead to higher detection accuracy, at the cost of efficiency. We expect this backward-query generation approach to hallucination detection can also be applied to planning tasks, where hallucinated plans will reverse-generate task descriptions that do not match well with the original goal.
C.3.2. Adversarial Prompting
  1. (3)
    Motivated by similar findings that language models are overfitting to QA datasets, Ramakrishna et al. (2023) present a novel framework for generating invalid questions to test the frequency of hallucinations output by new LLMs. The authors collect an augmented version of the DBpedia dataset (Lehmann et al., 2015) replaced with invalid questions to evaluate the hallucination rate of language models, and note that any deployment dataset could have been used instead. Specifically, Ramakrishna et al. manually create a list of $24$ question templates with tags for subjects and objects that can be filled in. A set of $100$ invalid questions using the templates is generated by sampling subjects and objects from disjoint facts in the original dataset, and ensuring the questions do not have valid answers in DBpedia. Additional invalid questions are produced by replacing dates within questions from TriviaQA (Joshi et al., 2017) with ones that do not exist. By passing each test question through various open-source and proprietary LLMs, the authors manually validate that each model has a tendency to hallucinate. Unfortunately, Ramakrishna et al. find that using automated evaluation metrics (discussed in Appendix D.1) do not correlate well with human annotations on their generated dataset, leaving room for future growth in aligned automatic evaluation metrics.
  2. (4)
    A more recent adversarial method from Uluoglakci and Temizel (2024) attempts to automatically generate hypothetical, invalid questions that language models should reject answering. For example, the hypothetical question, “What are the differences between Platypus LLM and Wolf LLM?” should be rejected since Wolf LLM is nonexistent, even though Platypus (Lee et al., 2023) is a real family of language models. Intuitively, if a model provides a plausible answer to the invalid question, the authors claim that either the fabricated term is not be in the model’s training set or the model has a higher tendency to hallucinate overall. As such, the authors propose a new framework to generate hypothetical questions, with which they create the HypoTermQA dataset. In particular, Uluoglakci and Temizel first query GPT-3.5 for $20$ popular topics and then $50$ hypothetical terms per topic, resulting in $790$ filtered fake phrases. To generate a diverse set of outputs, the model is set with a high temperature parameter. The authors argue that questions produced with only hypothetical terms will be easier to distinguish by language models, and thus, generate a set of similar terms with real meanings, in an attempt to deceive models. Similar, valid phrases were generated using three distinct approaches: querying GPT-3.5 directly, retrieving titles from Wikipedia with similar vector embeddings to the hypothetical terms, and taking the titles of Wikipedia passages whose definition embeddings are similar to those of the hypothetical terms. Finally, GPT-3.5 is instructed to generate plausible questions with filtered hypothetical and valid terms to produce a total of $19.5$K hypothetical and valid questions. Rather than relying on human annotation for identifying hallucinated responses, Uluoglakci and Temizel additionally propose HypoTermQA Score — the ratio of valid answers to the total number of hypothetical questions, automatically labeled by a proxy evaluator LLM agent (like methods in Section 4.3.3). In evaluation, the authors find that both open-source and proprietary models have over $90\%$ frequency of generating invalid responses. However, this frequency differed depending on the chosen evalautor model due to biases. Thus, proxy evaluator agents are also tested by comparing against manual annotation. Overall, the proposed dataset generation and automatic hallucination detection method show great promise in evaluating the factual hallucination tendency of language models. But, additional work should be done to handle biases in evaluator models, evaluate other characteristics of hallucinations, and consider other deployments outside QA.
C.3.3. Proxy Model
  1. (3)
@@ C.3.3. Proxy Model
    Similarly, Pacchiardi et al. (2024) develop a black-box lie detector for LLMs. In their case, the authors hypothesize that models that output a lie will produce different behaviors in future responses, like Azaria and Mitchell (2023). As such, at inference time, Pacchiardi et al. prompt the LLM with several binary questions (that may be completely unrelated to the original response) and collect yes/no answers. All the responses are concatenated into a single embedding that is input to the logistic regression model to predict the likelihood that the response was untruthful. To evaluate their lie detector, the authors assemble over $20$K questions from existing data sources on topics including general trivia (Vrandečić and Krötzsch, 2014; Welbl et al., 2017; Meng et al., 2022), basic arithmetic (Patel et al., 2021), common sense reasoning (Talmor et al., 2021), text translation (Tiedemann, 2012), and self-awareness (Perez et al., 2023). They additionally synthetically generate questions with unknowable answers as a control split (e.g., “What day is it?”). The authors find that the simple detector is mostly task- and model-agnostic once trained on a single dataset.
  2. (4)
    Chu et al. (2024) primarily rely on the LVLM GPT-4 to detect hallucinations within AI-generated videos. Video hallucinations are classified as static hallucinations, which occur in individual frames, or dynamic hallucinations, which occur across multiple frames. Here, the authors employ the generalized capabilities of GPT-4 to detect objects within keyframes of a video, summarize the video from the keyframes (which might be inconsistent with the original video generation prompt due to hallucinations), synthesize a temporal knowledge graph representing the changing relations among detected objects, detect inconsistencies along the knowledge graph, aggregate a hallucination score, and describe the detected hallucinations. Sora Detector is evaluated on a custom dataset collected by the authors, T2VHaluBench (described in Appendix D.2.4), and outperforms video hallucination detection ablation approaches that ignore knowledge graph construction, showcasing the usefulness of the representation. As such the structured approach shows promise in identifying undesired effects in generated videos, but the authors do not provide discussion on the possible hallucinations generated by GPT-4 when detecting and describing failures.
C.3.4. Grounding Knowledge
  1. (2)
@@ C.3.4. Grounding Knowledge
  2. (3)
    Another source of potential conflict that leads to hallucinations, is misalignment between a model’s capabilities and the user’s beliefs about what it can do. Zhang et al. (2024a) tackle this knowledge alignment problem and categorize alignment failures into four types:
@@
  3. (6)
    Lim and Shim (2024) tackle the issue of image generation models producing results that are factually noncompliant with the original prompt, by utilizing an external database to address inconsistencies. In particular, the authors confront three forms of image hallucinations: factual inconsistencies, outdated knowledge, and factual fabrications (unlikely generations). Given a prompt and an original (hallucinated) generation from a generator, the authors manually collect a ground-truth image describing the prompt using Google. An LVLM is employed to identify the differences between the retrieved and generated images, resulting in an instruction for how to edit the original generation to become factual. Finally, the original generation and editing instruction is passed through the generator to produce an updated result. Qualitatively, the pipeline produces more factually compliant and plausible generations. However, a human needs to manually choose the ground-truth image, the prompt correction generation model can hallucinate improper instructions, and the authors do not compare their method with other image hallucination mitigation methods.
  4. (7)
    Schramowski et al. (2023) present Safe Latent Diffusion (SLD), a method to provide additional guidance to image generation diffusion models at inference time to reduce the tendency of inappropriate generations. Intuitively, diffusion models iteratively remove noise from a generation until a plausible result is produced. Along the way however, inappropriate features like nudity or violence can be denoised, as seen in the training dataset. SLD updates the original prompt guidance encoding by computing the latent encoding of unsafe guidance, and shifting the original encoding away from the unsafe encoding given set hyperparameters. The resulting shifted “safe” guidance is far from unsafe embeddings, but still close enough to the original prompt guidance, to produce generations that attempt to follow the original prompt without inappropriate content. This augmented guidance is only provided after some warm-up steps of denoising to allow for the model to generate results close to the original prompt. The authors provide four sets of hyperparameters to configure the aggressiveness of the content filter. Using a model that detects inappropriate content, Schramowski et al. show that SLD has a lower probability of generating undesired features, and a user study supports that the safe generations are still preferred the same as (or more than) the unaugmented results. It is important to note that, while SLD combats undesired generations, there are still instances of inappropriate features present due to their representation in the training dataset. Furthermore, the proposed algorithm can be misused by shifting the augmented guidance toward the unsafe vector, producing more inappropriate features.
Appendix D Metrics and Evaluation Platforms